OpenStudy (anonymous):

differentiate f(x)=secx/1+tanx

OpenStudy (anonymous):

Is this what you're given?\[f(x)=\frac{ \sec(x) }{ 1+\tan(x) }\]

OpenStudy (anonymous):

What you have written would actually be\[f(x)=\sec(x)+\tan(x)\]

OpenStudy (anonymous):

its a quotient, thus commands application of quotion rule in my view

OpenStudy (anonymous):

I haven't taken a derivative, I'm merely trying to clarify what the original problem is.

OpenStudy (anonymous):

its a quotient, consider your first correction not the second

OpenStudy (anonymous):

Ah, okay, so just plug in your values for quotient rule where\[\frac{ f(x) }{ g(x) }=\frac{ f'(x)g(x)-f(x)g'(x) }{ (g(x))^{2} }\]

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

i bet brenar has a point there,

OpenStudy (loser66):

here What you have written would actually be f(x)=sec(x)+tan(x) that 's yours

OpenStudy (jhannybean):

\[\large \frac{d}{dx}[\frac{\sec(x)}{1+\tan(x)}]\]\[\large f'(x) = \frac{d}{dx}(\sec x) = \sec(x)\tan(x)\]\[\large g'(x) = \frac{d}{dx}[\tan(x)] =\sec^2(x)\]

OpenStudy (anonymous):

@Loser66 I wrote that because @falzedu had the wrong notation in his original equation, I merely wanted to clarify, because I knew what he had most likely intended. It's an order of operations thing.

OpenStudy (jhannybean):

Can you solve it now? lol

OpenStudy (jhannybean):

just use the quotient rule, Goodness.

OpenStudy (anonymous):

haha @Jhannybean

OpenStudy (loser66):

@Brenar got it, but I don't think there is something wrong with the question. quotient rule, done

OpenStudy (jhannybean):

\[\large \frac{[\sec(x)\tan(x)]*(1+\tan(x)) - \sec^2(x)[\sec(x)]}{[1+\tan(x)]^{2}}\]

OpenStudy (jhannybean):

That is ILLEGAL

OpenStudy (anonymous):

Because that's not what he wrote, he wrote f(x)=secx/1+tanx which is the same as\[f(x)=\sec(x)+\tan(x)\]

OpenStudy (anonymous):

Just... calm down and use quotient rule.

OpenStudy (jhannybean):

Stop being so knit-picky. He just forgot the parenthesis. Lol f(x) = (sec(x))/(1+tan(x))

OpenStudy (anonymous):

@Jhannybean I know, but I don't believe @Loser66 understood that.

terenzreignz (terenzreignz):

I always thought it was "nit-picky" But yeah, @Jhannybean 's right, the power of LaTeX wasn't used so pick the more probable question :D

OpenStudy (jhannybean):

lol KNITTING AND PICKING!

OpenStudy (loser66):

post a new post friend